OpenStudy (anonymous):

if f(x)=x+sq root 2-x and g(u)= u+2-u is it true that f=g?

OpenStudy (anonymous):

\[ f(x) = x+\sqrt{2}-x \]And \[ g(u) = u+2-u \]

OpenStudy (anonymous):

\[ f=g\iff f(x)-g(x)=0 \]

OpenStudy (anonymous):

\[ f(x)-g(x) = x+\sqrt{2}-x-(x+2-x) = \sqrt{2}-2 \neq 0 \]